OpenStudy (anonymous):

Help me out with this please? A drug trial is testing the effectiveness of two drugs. If 30 patients are given Drug A, 10 patients are given Drug B, and 90 patients are given a placebo, what is the probability that a patient will not be given a placebo ? A. 4/13 B. 3/13 C. 9/13 D. 1/13 Please give letter answer with response to why.

OpenStudy (anonymous):

Just find a common denominator and add them together, then simplify.

jagr2713 (jagr2713):

30 + 10 = 40 = 90+10+30 = 130 = 40/1130 = 4/13. So, answer is A

OpenStudy (anonymous):

Thanks.

jagr2713 (jagr2713):

Your're welcome